# A Framework for Situating Innovations, Opportunities, and Challenges in Advancing Vertical Systems with Large AI Models

### v1

@@ 5.1 Optimally situating and borrowing innovations
One of the key intended outcomes of the framework is to help situate the innovations such that they have increased potential for impact and also promote optimal use of resources. An important aspect of this is to consider when vertical-specific innovations, particularly on the modeling and interfacing front, could have broader impact across many verticals. For instance, studies across many verticals have found that adapting LLMs on benign datasets for healthcare, education, and law compromises their safety — making them more likely to respond to potentially harmful prompts [64]. Individual research teams across verticals therefore go through additional steps to first quantify the extent of compromise in safety upon fine-tuning and then, if the extent is unacceptable, improve the compromised safety of vertically-adapted LLMs [65, 66]. However, as opposed to investigating and addressing the underlying issues independently in many verticals, Peng et al. (2024) study the loss of safety in adapted LLMs—across several LLMs and datasets—to propose a metric for safety in LLM adaptation by visualizing its safety landscape [67]. The metric could be used across verticals to determine whether the LLM adaptation has compromised safety and the extent to which remedial strategies are required. This illustrates how addressing challenges that are encountered across several verticals as a vertical-agnostic properties could be more optimal in terms of impact and resource allocation.
@@
Similar examples exist in how vertical-specific interfacing challenges could point to broadly applicable problems that are optimally addressed as vertical-user intermediaries. As a specific example, the tendency of LLM-based chatbots to neglect crucial aspects of user interactions as they primarily focus on outcomes has been noted in many verticals, including their applications in therapeutic conversations [68], information seeking [69], writing code [70], or humanitarian frontline negotiations [71]. While vertical-specific strategies have been proposed (like inducing Gricean maxims to structure code-related interactions [70]), it is promising to assess whether such strategies generalize across verticals as a vertical-user intermediary. If such a challenge can be addressed in a vertical-agnostic manner, it could enable more effective vertical-user interactions across many verticals. It is worth emphasizing that situating innovations as vertical-agnostic properties (whether in modeling or interfacing) does not aim to discourage vertical-specific innovation as it is the latter that provides insights into what could work (snowballing effect). The objective here is to encourage researchers and practitioners to consider the potential impact of their innovation and aid others in adopting it to establish vertical-agnostic generalization.
@@
Conversely, optimality also requires that researchers and practitioners adopt vertical-agnostic strategies when they are effective in their specific verticals. For instance, while many vertical-specific works note that prompt engineering is often ad hoc, solutions like DSPy [72] provide a structured, vertical-agnostic approach to optimizing prompts, thereby addressing common challenges across multiple domains and saving resources.
5.2 Uncovering overlooked opportunities
@@
Guided by the philosophy of “turning frequent failures into signals”, the framework can also help uncover overlooked opportunities. A compelling example exists in the development and adoption of large language models for languages in South Asia. Individual teams have noted the limitations of LLMs trained on predominantly English data when they are used in non-English languages for applications in different verticals [73, 74, 75]. Such LLMs also lack the socio-cultural awareness to adapt to the needs of users in South Asian regions [76]. These failures have highlighted an opportunity for collaborative and participatory research to develop new LLMs for South Asian languages, such as BharatGPT [77] and SeaLLMs [78]. Since their deployment, these models have been adopted across many verticals by businesses, governments, and non-profits [79]. By identifying the pattern among the failures in vertical adoption and addressing the major issues at the large AI model-layer of the framework, the vertical-agnostic and vertical-specific layers can build on top of advanced off-the-shelf capabilities of the newer models. Such coordinated and targeted efforts also ensure that the burden of developing large AI models that are effective in specific verticals does not fall on individual teams as the process is restrictively resource intensive and carries environmental costs [80].
@@
Another bottleneck that exists in the vertical adoption of large models is careful handling of sensitive data – including health data [33], student data [34], and proprietary workflows [81]. Current large models readily allow adaptation via in-context learning or custom system instructions, which pose the risk of exposing sensitive data via jailbreak attacks [82]. However, Rajendran et al. (2024) [83] argue that cross-cohort cross-category integration — “the process of combining information from diverse datasets distributed across distinct, secure sites” — is important for adopting large models in verticals like healthcare. This provides an opportunity to develop large AI models or vertical-agnostic properties that facilitate adaptability while securely handling data without compromising on key performance metrics. Training large models for adaptability via approaches like meta-learning [84] or providing secure adaptability as a service [81] are some opportunities to accelerate vertical adoption.
@@
Beyond modeling-related opportunities, there are significant opportunities on the human-AI interaction and interface design fronts, both in vertical-specific as well vertical-agnostic settings. The large-scale user-adoption of AI is still relatively nascent [85, 86] and while modeling-related problems (both vertical-specific and vertical-agnostic) are being actively addressed by newer iterations of the large models, the interfacing of these capabilities with the intended users is still a major challenge. For instance, beyond chatbot-like interfaces, recent studies show structured media such as notebooks provide a flexible interface for incrementally creating and consuming information, which is also effective for clinically mandated documentation standards [87, 88, 89]. Additionally, micro-prompting and using interactive graphical objects (an interaction paradigm that can be applied to many verticals) has shown to enhance user satisfaction in human-AI interactions [90, 91, 92].
5.3 Communicating cross-disciplinary challenges
@@
By helping researchers situate their efforts within a common structure, the framework promotes streamlined collaboration among AI developers, vertical experts, and human-computer interaction researchers, ensuring that large models evolve into useful and trustworthy tools across a wide range of real-world applications. The framework also helps avoid the pitfalls of viewing large AI models as self-sufficient solutions [93, 22]. At the same time, it encourages vertical-specific teams to articulate their domain’s unique requirements and allows large-model developers to spot recurring problems that merit general-purpose fixes.
6 Discussion
@@
Our framework addresses the crucial gap between the remarkable capabilities of large AI models and the complex, context-specific needs that arise in real-world verticals. By modularizing the development pipeline into four layers — large AI models, vertical-agnostic properties, vertical adaptation, and vertical-user intermediaries — it offers a structured way to identify, situate, and address the challenges encountered when building practical AI systems. We also highlighted the dynamic interplay among these layers: insights from domain-specific challenges loop back into model-level improvements, and widespread interface challenges become generalizable solutions for human-AI interactions. To that end, the framework provides actionable guidance for effectively placing innovations, spotting opportunities that might otherwise be overlooked, and coordinating across disciplines to precipitate vertical-specific utility.
@@
Limitations and future directions: Our framework, while providing a valuable starting point for situating innovations, opportunities and challenges in vertical adoption of large AI models, has multiple limitations. The framework centers on end-users and their direct interactions with AI systems, leaving out broader social, institutional, and policy contexts. In the context of healthcare and well-being, De Choudhury et al. (2023) [94] adopt a Social Ecological Model to explore how AI can influence not just individuals but also caregiving institutions and society at large. While these wider sociotechnical considerations are crucial for fully understanding AI’s impact — especially in sensitive areas like mental health and telehealth — they fall beyond the scope of our current user-focused framework. Future work could integrate our framework with broader ecological models to capture the dynamism across multiple stakeholders and social layers, thereby offering a more holistic view of how AI systems both shape and are shaped by larger sociotechnical contexts. Furthermore, since our framework aims to cover the development pipeline that transforms large AI models into vertical systems, it does not dive deep into individual challenges. For instance, prior studies have developed frameworks around specific aspects like robustness, interpretability, explainability, and human-AI interaction and communication [95, 96, 97]. Future work could involve a large-scale study that balances high-level coverage of the entire pipeline with deeper dives into each sub-issue. This would require a multi-disciplinary collaboration of experts in those sub-issues, ensuring that the final systems remain grounded in real user needs while thoroughly addressing the nuanced challenges in each aspect of deployment.

### v2

@@ A Framework for Advancing Vertical Systems With AI Models
One of the key intended outcomes of the framework is to help situate the innovations such that they have increased potential for impact and also promote optimal use of resources. An important aspect of this is to consider when vertical-specific innovations, particularly on the modeling and interfacing front, could have broader impact across many verticals. For instance, studies across many verticals have found that adapting LLMs on benign datasets for healthcare, education, and law compromises their safety — making them more likely to respond to potentially harmful prompts (Qi et al. 2024). Individual research teams across verticals therefore go through additional steps to first quantify the extent of compromise in safety upon fine-tuning and then, if the extent is unacceptable, improve the compromised safety of vertically-adapted LLMs (Niknazar et al. 2024; De Freitas et al. 2024). However, as opposed to investigating and addressing the underlying issues independently in many verticals, Peng et al. (2024) study the loss of safety in adapted LLMs—across several LLMs and datasets—to propose a metric for safety in LLM adaptation by visualizing its safety landscape (Peng et al. 2024). The metric could be used across verticals to determine whether the LLM adaptation has compromised safety and the extent to which remedial strategies are required. This illustrates how addressing challenges that are encountered across several verticals as a vertical-agnostic properties could be more optimal in terms of impact and resource allocation.
@@
Similar examples exist in how vertical-specific interfacing challenges could point to broadly applicable problems that are optimally addressed as vertical-user intermediaries. As a specific example, the tendency of LLM-based chatbots to neglect crucial aspects of user interactions as they primarily focus on outcomes has been noted in many verticals, including their applications in therapeutic conversations (Zhou et al. 2024), information seeking (Sharma, Liao, and Xiao 2024), writing code (Bajpai et al. 2024), or humanitarian frontline negotiations (Ma et al. 2024). While vertical-specific strategies have been proposed (like inducing Gricean maxims to structure code-related interactions (Bajpai et al. 2024)), it is promising to assess whether such strategies generalize across verticals as a vertical-user intermediary. If such a challenge can be addressed in a vertical-agnostic manner, it could enable more effective vertical-user interactions across many verticals. It is worth emphasizing that situating innovations as vertical-agnostic properties (whether in modeling or interfacing) does not aim to discourage vertical-specific innovation as it is the latter that provides insights into what could work (snowballing effect). The objective here is to encourage researchers and practitioners to consider the potential impact of their innovation and aid others in adopting it to establish vertical-agnostic generalization.
@@
Conversely, optimality also requires that researchers and practitioners adopt vertical-agnostic strategies when they are effective in their specific verticals. For instance, while many vertical-specific works note that prompt engineering is often ad hoc, solutions like DSPy (Khattab et al. 2023) provide a structured, vertical-agnostic approach to optimizing prompts, thereby addressing common challenges across multiple domains and saving resources.
Uncovering overlooked opportunities
@@
Guided by the philosophy of “turning frequent failures into signals”, the framework can also help uncover overlooked opportunities. A compelling example exists in the development and adoption of large language models for languages in South Asia. Individual teams have noted the limitations of LLMs trained on predominantly English data when they are used in non-English languages for applications in different verticals (Jin et al. 2024; Kumar et al. 2024; Verma et al. 2022a). Such LLMs also lack the socio-cultural awareness to adapt to the needs of users in South Asian regions (Pawar et al. 2024). These failures have highlighted an opportunity for collaborative and participatory research to develop new LLMs for South Asian languages, such as BharatGPT (BharatGPT 2025) and SeaLLMs (Nguyen et al. 2023). Since their deployment, these models have been adopted across many verticals by businesses, governments, and non-profits (CoRover AI ). By identifying the pattern among the failures in vertical adoption and addressing the major issues at the large AI model-layer of the framework, the vertical-agnostic and vertical-specific layers can build on top of advanced off-the-shelf capabilities of the newer models. Such coordinated and targeted efforts also ensure that the burden of developing large AI models that are effective in specific verticals does not fall on individual teams as the process is restrictively resource intensive and carries environmental costs (Strubell, Ganesh, and McCallum 2020).
@@
Another bottleneck that exists in the vertical adoption of large models is careful handling of sensitive data – including health data (Pan et al. 2024), student data (Yang and Beil 2024), and proprietary workflows (Tang et al. 2024). Current large models readily allow adaptation via in-context learning or custom system instructions, which pose the risk of exposing sensitive data via jailbreak attacks (Liu et al. 2023). However, Rajendran et al. (2024) (Rajendran et al. 2024) argue that cross-cohort cross-category integration — “the process of combining information from diverse datasets distributed across distinct, secure sites” — is important for adopting large models in verticals like healthcare. This provides an opportunity to develop large AI models or vertical-agnostic properties that facilitate adaptability while securely handling data without compromising on key performance metrics. Training large models for adaptability via approaches like meta-learning (Verma et al. 2024b) or providing secure adaptability as a service (Tang et al. 2024) are some opportunities to accelerate vertical adoption of large AI models.
@@
Beyond modeling-related opportunities, there are significant opportunities on the human-AI interaction and interface design fronts, both in vertical-specific as well vertical-agnostic settings. The large-scale user-adoption of AI is still relatively nascent (Houter 2024; Abril 2024) and while modeling-related problems (both vertical-specific and vertical-agnostic) are being actively addressed by newer iterations of the large models, the interfacing of these capabilities with the intended users is still a major challenge. For instance, beyond chatbot-like interfaces, recent studies show structured media such as notebooks provide a flexible interface for incrementally creating and consuming information, which is also effective for clinically mandated documentation standards (Cheng et al. 2024; Wang, Dai, and Edwards 2022; Adler-Milstein et al. 2022). Additionally, micro-prompting and using interactive graphical objects (an interaction paradigm that can be applied to many verticals) has shown to enhance user satisfaction in interactions between human and AI systems (Suh et al. 2023; Jiang et al. 2023; Butler et al. 2024).
Communicating cross-disciplinary challenges
@@
By helping researchers situate their efforts within a common structure, the framework promotes streamlined collaboration among AI developers, vertical experts, and human-computer interaction researchers, ensuring that large models evolve into useful and trustworthy tools across a wide range of real-world applications. The framework also helps avoid the pitfalls of viewing large AI models as self-sufficient solutions (Blodgett and Madaio 2021; Bommasani et al. 2021). At the same time, it encourages vertical-specific teams to articulate their domain’s unique requirements and allows large-model developers to spot recurring problems that merit general-purpose fixes.
Discussion
Positioning with respect to other frameworks: As mentioned earlier, most existing frameworks focus on specific facets of adopting large AI models into verticals. For instance,  Ehsan et al. (2023) chart the sociotechnical gap in explainable artificial intelligence, presenting a framework that conceptualizes the gap between model outputs and human interpretability needs. Similarly,  Goldstein and Sastry (2024) present a framework for estimating the malicious use of systems built with advanced AI models. In the domain of conversational search, Azzopardi et al. (2024) develop a conceptual model of agent–user interaction, mapping out the dialogue acts and decision points that drive information-seeking conversations. Moving beyond facets like explainability and interfacing, at a systems level, Yan et al. (2024) provide a list of engineering-focused best-practices for LLM-based applications, emphasizing components like rigorous evaluation, retrieval-augmented generation, fine-tuning, and guardrails to integrate large models into products. In the same vein, institutions have also put forth adoption guides: a U.S. Department of Energy report ( 2022) underscores the need for robust data pipelines, high-performance computing, and reproducible model infrastructures to accelerate AI uptake in critical domains. Likewise, the NIST AI Risk Management Framework ( NIST) focuses on processes for identifying and mitigating AI system risks, defining functions such as Govern, Map, Measure, and Manage to ensure trustworthy development and deployment. Industry frameworks like IBM’s “AI Ladder” ( 2025) similarly prescribe sequential steps (e.g., modernize, collect, organize, analyze, infuse) to guide organizations in scaling AI from data preparation to integration into workflows.
Each of these works offers valuable conceptual frameworks, but notably each isolates a specific sub-problem or component of AI adoption into verticals — be it explainability, user interfacing, engineering best practices, infrastructure needs, or product management cycles — rather than providing a unified framework. In contrast, our layered framework offers a holistic and modular structure that ties these aspects together. It spans from the core large-model layer, through vertical-agnostic properties and vertical-specific adaptation, up to user-facing intermediaries, explicitly situating innovations and challenges in context. Within this integrated view, we present several perspectives that could enable an ecosystem of efficient and effective vertical adoption of AI models — e.g., how improvements in one layer (for example, a robustness technique at the model layer or a novel interfacing paradigm at the user-intermediary layer) can propagate benefits across other layers and across domains. Cross-cutting concerns, such as data privacy, fairness, or trust, are better addressed at right layers such that they are not repeatedly “solved” in silos.
By providing a shared vocabulary and clear abstraction boundaries, the framework enables researchers and practitioners to communicate and build upon each other’s advances. Our framework not only advocates to reduce redundant effort (teams in different verticals can reuse solutions or insights from analogous layers) but also establishes a basis for diving deeper into aspect-specific frameworks as needed. For instance, an explainability framework (as in the framework by Ehsan et al.) slots naturally into our vertical-user intermediary layer, and a conversational interaction model (as in Azzopardi et al.’s framework) can be viewed as a vertical-specific interface component. In essence, the layered approach “situates” specialized innovations within a bigger picture. All of these qualities address the unmet need for an integrated conceptual models for AI adoption: one that modularly encompasses the full pipeline of transforming a large AI model into a domain-specific, user-facing system, and thereby complements prior frameworks.
Adaptability with the changing interaction paradigm: While our framework primarily focuses on human-AI interactions, it is inherently adaptable to emerging ways of interactions with large models. For instance, Model Context Protocol (MCP) is an open standard recently designed to facilitate interactions between large models and external tools (Anthropic 2024). As agentic-AI workflows gain prominence, understanding signals and feedback from model-tool interactions becomes crucial for further improvements. However, such signals could differ from traditional measures commonly used in human-AI interaction paradigm (such as user trust) and may include newer measures such as agentic-coordination capabilities, and tool-usage efficiency. Our framework provides the agility to incorporate such changes. For example, in this case, an additional layer parallel to the vertical-user intermediary layer could be introduced which specifically adapted to analyze multi-agent system interactions.
Limitations and future directions: Our framework, while providing a valuable starting point for situating innovations, opportunities and challenges in vertical adoption of large AI models, has multiple limitations. The framework centers on end-users and their direct interactions with AI systems, leaving out broader social, institutional, and policy contexts. In the context of healthcare and well-being, De Choudhury et al. (2023) (De Choudhury, Pendse, and Kumar 2023) adopt a Social Ecological Model to explore how AI can influence not just individuals but also caregiving institutions and society at large. While these wider sociotechnical considerations are crucial for fully understanding AI’s impact — especially in sensitive areas like mental health and telehealth — they fall beyond the scope of our current user-focused framework. Future work could integrate our framework with broader ecological models to capture the dynamism across multiple stakeholders and social layers, thereby offering a more holistic view of how AI systems both shape and are shaped by larger sociotechnical contexts. Furthermore, since our framework aims to cover the development pipeline that transforms large AI models into vertical systems, it does not dive deep into individual challenges. For instance, prior studies have developed frameworks around specific aspects like robustness, interpretability, explainability, and human-AI interaction and communication (Li et al. 2023; Fragiadakis et al. 2024; Bansal et al. 2024). Future work could involve a large-scale study that balances high-level coverage of the entire pipeline with deeper dives into each sub-issue. This would require a multi-disciplinary collaboration of experts in those sub-issues, ensuring that the final systems remain grounded in real user needs while thoroughly addressing the nuanced challenges in each aspect of deployment.
Conclusion
@@
Our framework addresses the crucial gap between the remarkable capabilities of large AI models and the complex, context-specific needs that arise in real-world verticals. By modularizing the development pipeline into four layers — large AI models, vertical-agnostic properties, vertical adaptation, and vertical-user intermediaries, our framework offers a structured way to identify, situate, and address the challenges encountered when building practical AI systems. We also highlighted the dynamic interplay among these layers: insights from domain-specific challenges loop back into model-level improvements, and widespread interface challenges become generalizable solutions for human-AI interactions. To that end, the framework provides actionable guidance for effectively placing innovations, spotting opportunities that might otherwise be overlooked, and coordinating across disciplines to precipitate vertical-specific utility.
@@
Acknowledgments
This research is based upon work supported in part by NSF grants CNS-2154118, ITE2137724, ITE-2230692, CNS2239879, NIH grants R01 MH117172 and P50 MH115838, Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112290102 (subcontract No. PO70745), CDC, and funding from Microsoft and the American Foundation for Suicide Prevention (AFSP). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the position or policy of NSF, NIH, DARPA, DoD, SRI International, CDC, Microsoft, or AFSP, and no official endorsement should be inferred. Gaurav is partly supported by the JPMorgan AI Research PhD Fellowship and the Snap Research Fellowship.